# Soteria: Language-Specific Functional Parameter Steering for Multilingual Safety Alignment

### v1

@@ 5.4 Baselines
(1) Safety arithmetic Hazra et al. (2024a): The safety arithmetic framework improves model safety across the base, fine-tuned, and edited models where safety risks emerge due to fine-tuning artefacts, or unintended editing consequences. It adjusts model parameters and realigns the latent space to reduce harmful outputs and ensure safer content generation.
(2) Resta Bhardwaj et al. (2024): This method restores safety in fine-tuned LLMs by adding a safety vector from the difference between a safety-aligned and an unaligned model. It counteracts safety degradation from supervised fine-tuning and enhances alignment using drop and rescale (DARE) Yu et al. (2024) to remove redundant delta parameters before applying Resta.
@@
(3) Self-defense Deng et al. (2024b): We could not compare the self-defense method which suggests that simple fine-tuning with a specific dataset can restore multilingual safety, due to the unavailability of the dataset mentioned in the paper.
Figure 5: Comparison of Soteria with other baselines [^7] We define average of High resources as High, and similarly for Mid and Low. This also holds for Figure 6 and Table 2..
@@ 6 Main results
General capabilities: We evaluate our framework’s impact on overall model capabilities using utility tests (MMLU Hendrycks et al. (2021) 5-shot and TruthfulQA Lin et al. (2022)). The results closely mirror each base model’s performance. For the safe version of Llama 3.1, we observe the MMLU performance at 72.9 (vs. 73 from the baseline), and TruthfulQA at 44.14 (vs. 44.14 for the baseline). The safe version of Qwen exactly matched its base values (70.3, 54.2). Mistral yielded 61.79 MMLU (vs. 61.84) and 59.34 TruthfulQA (vs. 59.37), while Phi also retained its baseline scores of 69 (MMLU) and 64 (TruthfulQA).
Comparison with the baselines: We compare Soteria with three English-centric safety alignment methods as discussed above – safety-arithmetic, Resta, and TIES – by examining the ASR values for high-, medium-, and low-resource languages. Figure 7 presents the results for two models, Llama 3.1 and Qwen 2, using the Multijail and XThreatBench datasets. Across all baselines, Soteria consistently achieves the lowest ASR. On Llama 3.1 with the Multijail dataset, the baseline method’s ASR ranges from 30–40% in high-resource languages, while for Soteria it is about 15–20%. Both TIES and Resta provide moderate decreases (30–35%), and safety-arithmetic does slightly better (25–30%). However, Soteria consistently outperforms these methods by 5–10%. Similar trends hold for medium- and low-resource languages. A comparable trend is also observed from Qwen 2. For Multijail, the baseline ASR is approximately 28–30% in high-resource settings, whereas TIES, Resta, and safety-arithmetic reduce it to 20–25%. Soteria pushes the ASR even lower, to around 15–20%. These findings also generalize to XThreatBench, reinforcing the robustness of Soteria across diversely resourced languages, models and datasets.
@@ 7 Language universals
Figure 6: Trade-off between ASR and % heads probed.
8 LLM jailbreaks
@@
10 Conclusion
11 Limitation
A key limitation of Soteria lies in its reliance on per-language functional neuron identification, which requires accurate language segmentation and task-based data in each target language. In practice, resource constraints, limited training data, and complexities in script variation or morphology can reduce the precision of head selection. Moreover, although Soteria improves safety across many languages, it does not guarantee comprehensive coverage of every cultural nuance or emergent harmful behavior.
12 Ethical Consideration
Appendix A Additional experiment
@@
A.1 Result for XSafety dataset
@@
A.2 XSafety (language universal)
@@
Appendix B Attention head patterns and their implications
@@
One intriguing characteristic of LLMs is how their top-valued language‐specific attention heads tend to cluster by resource level of the language. Analyses of a smaller-parameter model (e.g., Llama 3.1 8B‐parameter variant) reveal that high‐resource languages (such as English, Chinese, Spanish, German, and French) and mid‐resource languages (such as Hindi, Arabic, Thai, and Bulgarian) exhibit peak attention heads in roughly the same mid‐level layers (e.g., layers 12–20 with head indices 16–24). Meanwhile, for low‐resource languages the strongest attention heads manifest in later layers (e.g., layers 28–31 with head indices 15–23) (see Figure 7).
@@
(1) Language-specific universal heads: Despite the differences in where each language’s top heads appear, some heads consistently contribute to cross‐lingual understanding – the so‐called “universal” heads. Identifying and enhancing these universal heads can make the model’s latent space more cohesive across languages, improving zero‐shot or few‐shot performance for underrepresented languages.
@@
(2) Future directions: Beyond raw performance, attention‐head analysis also provides new insights to tackle task-specific attention heads, misalignment, and hallucination issues. If certain heads consistently carry problematic correlations, shifting or refining their latent space (“steer them to a safe side”) can enhance overall alignment and trustworthiness.
@@
These findings underscore the delicate interplay between multilingualism and architectural depth in multilingual models. By homing in on the most influential heads and understanding why they appear where they do, we gain powerful levers for improving cross‐lingual performance, minimizing unsafe content generation, and facilitating more robust language support, even for the world’s most resource sparse tongues.

### v2

@@ 5.1 Baselines
(1) Safety arithmetic Hazra et al. (2024b): The safety arithmetic framework improves safety across the base, fine-tuned, and edited models where safety risks emerge due to fine-tuning artefacts, or unintended editing consequences. It adjusts the parameters and realigns the latent space to reduce harmful outputs and ensures safer content generation.
(2) Resta Bhardwaj et al. (2024): It restores safety in fine-tuned LLMs by adding a safety vector equal to the difference between a safety-aligned and an unaligned model. It further enhances alignment using drop and rescale (DARE) Yu et al. (2024) to remove redundant delta parameters before applying Resta.
@@
(4) Self-defense Deng et al. (2024b): We could not compare the self-defense method, which suggests that simple fine-tuning with a specific dataset can restore multilingual safety, due to the unavailability of the dataset mentioned in the paper.
77footnotetext: We define average of High resources as High, and similarly for Mid and Low. This also holds for Figure 7 and Table 3.
@@ 6 Main results
Comparison with the baselines: We compare Soteria with three English-centric safety alignment methods as discussed above – safety-arithmetic, Resta, and TIES – by examining the ASR values for high-, medium-, and low-resource languages (also see Appendix for performance of Soteria with random attention heads). Figure 6 presents the results for two models, Llama 3.1 and Qwen 2, using the Multijail and XThreatBench datasets. Across all baselines, Soteria consistently achieves the lowest ASR. On Llama 3.1 with the Multijail dataset, the baseline method’s ASR ranges from 30–40% in high-resource languages, while for Soteria it is about 15–20%. Both TIES and Resta provide moderate decreases (30–35%), and safety-arithmetic does slightly better (25–30%). However, Soteria consistently outperforms these methods by 5–10%. Similar trends hold for medium- and low-resource languages. A comparable trend is also observed from Qwen 2. For Multijail, the baseline ASR is approximately 28–30% in high-resource settings, whereas TIES, Resta, and safety-arithmetic reduce it to 20–25%. Soteria pushes the ASR even lower, to around 15–20%. These findings also generalize to XThreatBench, reinforcing the robustness of Soteria across diversely resourced languages, models and datasets.
Figure 5: Comparison of harmful and safe responses generated using Qwen model across multilingual examples. Panel (a) illustrates harmful outputs produced by the Safety Arithmetic (SA) method, highlighting sub-sentences annotated with sequence attribution scores indicating their contribution to harmfulness. Panel (b) demonstrates safe responses produced using our proposed method, Soteria, with sub-sentence scores reflecting improved safety. Examples include texts in Bengali and Hindi.
@@ 7 Language universals
8 Interpreting via attribution maps
Figure 6: Comparison of Soteria with other baselines99footnotemark: 9.
To enhance interpretability, we conduct an additional analysis using Captum [^10] https://captum.ai/, comparing input attributions for harmful questions and their corresponding answers. We observe that SA (our closest competitor) often generates harmful responses in multilingual settings due to diffused attribution across input tokens. In contrast, Soteria, assigns higher attribution to key harmful tokens, enabling more reliable refusal behaviour. This demonstrates Soteria’s superior capability in producing safer responses across languages. In Figure 5, the scores associated with sub-sentences represent sequence attribution scores, scaled between 0 and 1. Here, a score of 0 indicates no attribution, while a score approaching 1 signifies very high attribution. Higher attribution scores demonstrate the importance of the corresponding sequence within the input toward generating the given output. Indirectly, these are the words mostly attended to by the model. The sequence of words with a deeper color gradient shows that they have a higher impact on the generated output. In the Bengali example, for instance, the sequence images of child abuse is correctly identified as harmful by Soteria’s attention mechanism, while this is not the case for SA. Since Soteria can track harmful instances in the input more correctly, it can better understand and generate an ethical output. The same observations hold for other instances and languages.
9 Conclusion
10 Limitation
A key limitation of Soteria lies in its reliance on per-language functional neuron identification, which requires accurate language segmentation and task-based data in each target language. In practice, resource constraints, limited training data, and complexities in script variation or morphology can reduce the precision of head selection. Moreover, although Soteria improves safety across many languages, it does not guarantee comprehensive coverage of every cultural nuance or emergent harmful behaviour.
11 Ethical consideration
Appendix A General capabilities
We evaluate our framework’s impact on overall model capabilities using utility tests MMLU Hendrycks et al. (2021) 5-shot and TruthfulQA Lin et al. (2022) (see Table 2). The results closely mirror each base model’s performance. For the safe version of Llama 3.1, we observe the MMLU performance at 72.9 (vs. 73 from the baseline), and TruthfulQA at 44.14 (vs. 44.14 for the baseline). The safe version of Qwen exactly matched its base values (70.3, 54.2). Mistral yielded 61.79 MMLU (vs. 61.84) and 59.34 TruthfulQA (vs. 59.37), while Phi also retained its baseline scores of 69 (MMLU) and 64 (TruthfulQA).
Model MMLU (5-shot) TruthfulQA
Safe Baseline Safe Baseline
Llama 3.1 72.9 73.0 44.14 44.14
Qwen 70.3 70.3 54.2 54.2
Mistral 61.79 61.84 59.34 59.37
Phi 69.0 69.0 64.0 64.0
Table 2: Performance comparison of safe and baseline versions across utility benchmarks: MMLU (5-shot) and TruthfulQA.
Appendix B Evaluation prompt
Following Banerjee et al. (2024b) we use below prompt for safety evaluation.
Given the following text, analyze and assess its ethical implications. Determine whether the content adheres to ethical standards or if there are elements that raise ethical concerns.
Appendix C LLM jailbreaks
@@
Figure 7: Trade-off between ASR and % heads probed.
Appendix E Additional experiment
@@
E.1 Result for XSafety dataset
@@
Figure 9: Examples of culturally grounded AG News entries and their translations into Bengali, Spanish, French, and German. Despite linguistic diversity, key cultural references remain intact, supporting faithful cross-lingual task evaluation.
@@
E.2 XSafety (Language Universal)
@@
Appendix F Attention head patterns and their implications
@@
One intriguing characteristic of LLMs is how their top-valued language-specific attention heads tend to cluster by resource level of the language. Analyses of a smaller-parameter model (e.g., Llama 3.1 8B-parameter variant) reveal that high-resource languages (such as English, Chinese, Spanish, German, and French) and mid-resource languages (such as Hindi, Arabic, Thai, and Bulgarian) exhibit peak attention heads in roughly the same mid-level layers (e.g., layers 12–20 with head indices 16–24). Meanwhile, for low-resource languages the strongest attention heads manifest in later layers (e.g., layers 28–31 with head indices 15–23) (see Figure 8).
@@
(1) Language-specific universal heads: Despite the differences in where each language’s top heads appear, some heads consistently contribute to cross-lingual understanding – the so-called “universal” heads. Identifying and enhancing these universal heads can make the model’s latent space more cohesive across languages, improving zero-shot or few-shot performance for underrepresented languages.
@@
(2) Future directions: Beyond raw performance, attention-head analysis also provides new insights to tackle task-specific attention heads, misalignment, and hallucination issues. If certain heads consistently carry problematic correlations, shifting or refining their latent space (“steer them to a safe side”) can enhance overall alignment and trustworthiness.
@@
These findings underscore the delicate interplay between multilingualism and architectural depth in multilingual models. By homing in on the most influential heads and understanding why they appear where they do, we gain powerful levers for improving cross-lingual performance, minimizing unsafe content generation, and facilitating more robust language support, even for the world’s most resource sparse tongues.
Appendix G Sample corrupted prompts
(a) Sample from Adult Content category.
(b) Sample from Child Abuse Content category.
Figure 10: Examples from two sensitive content categories: (a) Adult content, and (b) Child abuse content, extracted as part of our dataset curation.
For the corrupted prompt, we set the prompt in such a way that each input is matched with a random output (see Table 11). We follow the same prompt corruption technique given in Todd et al. (2024).
Figure 11: Example of corrupted prompts.
Appendix H Cultural fidelity in translated task datasets
Resource level ASR MMLU % Heads
High Resource 0.31 72.9 25%
0.20 72.9 50%
0.18 72.8 75%
0.17 72.6 100%
Mid Resource 0.35 72.9 25%
0.23 72.9 50%
0.22 72.9 75%
0.19 72.7 100%
Low Resource 0.35 72.9 25%
0.28 72.9 50%
0.29 72.8 75%
0.29 72.8 100%
Table 6: ASR and MMLU scores by % heads retained across different resource levels.
When constructing multilingual task datasets by translating English inputs (e.g., AG News, sentiment analysis) into target languages, there is a potential concern that culturally sensitive references may not be accurately preserved, particularly in low-resource languages. To investigate this, we conduct a qualitative assessment of translated inputs across multiple languages, examining whether core cultural entities and contexts remain semantically aligned with the original.
Figure 9 presents examples from the AG News dataset, including instances that mention religious headwear, ethnonational identities, and interfaith holidays. These examples are translated into Bengali, Spanish, French, and German. The translations preserve high-fidelity references to key cultural elements, such as “headscarf”, “Irish-Moroccan”, “Christmas”, “Hanukkah”, and “Kwanzaa”. We observe that key semantic cues are retained even in low-resource languages like Bengali, thereby allowing meaningful category predictions to be made post-translation.
Appendix I Ablation: Random attention head selection
To further understand the efficacy of Soteria, we conduct an ablation experiment where attention heads were randomly selected rather than identified via our causal analysis.
We observe that while random selection yields some improvements over the base model, it is consistently inferior to Soteria across both MultiJail and XThreatBench datasets. This reaffirms the importance of our language-specific functional head identification strategy. Detailed ASR values across different resource categories and models are presented in Table 7.
Dataset Model High Mid Low
Qwen LLaMA Qwen LLaMA Qwen LLaMA
MultiJail Base 0.24 0.42 0.22 0.44 0.30 0.39
Random 0.21 0.32 0.20 0.29 0.29 0.34
Soteria (ours) 0.11 0.19 0.11 0.24 0.24 0.23
XThreatBench Base 0.12 0.21 0.15 0.25 0.16 0.20
Random 0.10 0.19 0.12 0.23 0.16 0.18
Soteria (ours) 0.07 0.12 0.12 0.18 0.16 0.16
Table 7: ASR comparison across Base, Random Attention Head Selection, and Soteria methods for two benchmark datasets (lower is better).
Appendix J Hyperparameter details
J.1 Key hyperparameters
Our framework introduces two main hyperparameters:
Parameter percentage: We restrict updates to only 3% of model parameters, specifically the O-projection weights associated with identified functional heads. This low-rank intervention significantly reduces harmful outputs without degrading general utility. We conduct an ablation study to quantitatively assess the impact of our modifications on the general capabilities of the model. To empirically validate that the influence on the model’s overall performance is minimal, we evaluate the modified model on the standard MMLU benchmark. The experimental results, presented in Table 6, confirm that these sparse parameter modifications have a negligible effect on the model’s general performance.
Lambda ($\lambda$): A scaling factor applied to the harm vector ($\hat{H}_{v}$) during safety steering (see Equation 7). We empirically found $\lambda\in[1,2]$ to be effective.
J.2 Fine-tuning Configuration for the Harmful Model
We fine-tune the harmful model using default configurations from the LLaMA Factory [^13] https://llamafactory.readthedocs.io/en/latest/ framework. The fine-tuning dataset comprises harmful queries and responses. The exact hyperparameter values are summarized in Table 12.
Figure 12: Identified top 20 heads for Llama 3.1 8B for all languages.